OpenStudy (anonymous):

x=7-y x y 0 is what 2 is what 7 is what

OpenStudy (anonymous):

x=0 , y=7 x=2 , y=5 x=7 ,y=0

OpenStudy (anonymous):

Just plug in the values of x in that equation to get the value of y.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

Well if you want to find y at x=0 , put x=0 in the equation x=7-y. You get 0=7-y. So y=7.

OpenStudy (anonymous):

what does 0.2.7=

OpenStudy (anonymous):

0.2.7? What do you mean?

OpenStudy (anonymous):

how do i get the answers for 0,2,7 with y what does y = in each of them

OpenStudy (anonymous):

y = 7,5,0 .

OpenStudy (anonymous):

y=what for7, y= what for 2. y= what for 0

OpenStudy (turingtest):

x=7-y x y 0 is what to solve this put the 0 where we have an x in the equation to get 0=7-y can you solve this for y ?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

would it be 7

OpenStudy (anonymous):

I am sorry it is 0

OpenStudy (turingtest):

it would be 7

OpenStudy (turingtest):

add y to both sides and you get y=7

OpenStudy (turingtest):

0=7-y +y +y y=7

OpenStudy (turingtest):

x=7-y x y 2 is what to answer this do the same thing as last time, but put a 2 where the x was

OpenStudy (anonymous):

so for 2 do i 7-2 for 2

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

2=7-y solve for y